Derive the constant A for a normalized one-dimensional Maxwellian distribution with variation from – infinity to + infinity.
F(u)= Ae-mu2/kT
1
Expert's answer
2020-05-12T09:32:02-0400
"a=\frac{m}{kT}"

"F(u)=Ae^{-0.5au^2}"

"\int_{-\infty}^{\infty}Ae^{-0.5au^2}du=1"

"\int_{-\infty}^{\infty}e^{-0.5au^2}du=\sqrt{\frac{2\pi}{a}}"

Thus,


"A=\sqrt{\frac{m}{2\pi kT}}"